OpenStudy (anonymous):

Consider the function f(x) = (1/2)x - 1 Find the inverse.

OpenStudy (anonymous):

Please if anyone can help

OpenStudy (nikita2):

y = f(x) so invers is (y+1)*2

OpenStudy (anonymous):

my book says y=f^-2(x)=2x+2 =o

OpenStudy (anonymous):

f^-1 i mean ^^

OpenStudy (anonymous):

\[f(x)=x/2-1\] where \[f(x)=y\] so \[2(y-1)=x\] replace all x with y, the inverse function is: \[y=2(x-1)\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

np :)